OpenStudy (anonymous):

The solution of the equation (x+1)/4-(2x-1)/6=x is? Help! lol

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

\[\frac{12(x+1)}{4}-\frac{12(2x-1)}{6}=12x\] i multiplied everything by 12 \[3(x+1)-2(2x-1)=12x\]

OpenStudy (anonymous):

idk wat have you tried so far

myininaya (myininaya):

3x+3-4x+2=12x -x+5=12x 5=12x+x 5=13x x=5/13

OpenStudy (annon):

(6(x+1)-(4(2x-1))/24 = x

OpenStudy (annon):

get the 24 over by multiplyiing both sides by 24... (6(x+1)-(4(2x-1))=24x

OpenStudy (annon):

distribute... 6x+6-8x+4=24x

myininaya (myininaya):

x=10/26=5/13

OpenStudy (annon):

oops ya add

OpenStudy (anonymous):

myininya got it right xD. Umm how did you get -4x+2 i got -4x-2

OpenStudy (annon):

so... -2x+10=24x then 10=24x+2x and x=10/26

OpenStudy (anonymous):

myininya got it right xD. Umm how did you get -4x+2 i got -4x-2

OpenStudy (annon):

distrbute negative sign

OpenStudy (anonymous):

Ohh damn i need to remember its a fraction lol

OpenStudy (anonymous):

:1